EDITORIAL

Terry Goodman

It is with a great sense of loss and sadness that we inform you of the death of Dr. Vince Edmondson. Vince passed away in June after a lengthy battle with brain cancer. Vince had served for a number of years as our Associate Editor for Computer Science and Applied Mathematics. While we will be able to replace him as Associate Editor, we will not be able to replace him as our colleague and friend. He will be missed greatly by our department and we are thankful for his friendship and enthusiastic support of MJMS.

Once again we want to use our editorial space for this issue to thank and acknowledge those individuals who serve as referees for the MJMS. We believe one of the strengths of our journal is the use of a thorough and consistent review process. The persons listed below play a vital role in that review process and we appreciate and applaud their efforts. Any reader who is interested in serving as a referee for the MJMS should complete the Referee Applicaton Form found after the Announcements.

List of Referees

Mangho Ahuja, Cape Girardeau, Missouri Hadi Y. Alkahby, Dillard University Jamuna P. Ambasht, Benedict College Mohammad K. Azarian, University of Evansville Seung–Jin Bang, Seoul, Korea Miron Bekker, University of Missouri–Rolla Gerald Bergum, Brookings, South Dakota John E. Bernard, University of Texas–Pan American David A. Brown, Ithaca College Jim Bruening, Southeast Missouri State University Larry Cammack, Central Missouri State University G. Alan Cannon, Southeastern Louisiana University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Eungchun Cho, Kentucky State University Steve Clark, University of Missouri–Rolla Gregory Davis, University of Wisconsin–Green Bay Bryan Dawson, Union University

List of Referees (cont.)

Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Albert Dixon, College of the Ozarks Clayton Dodge, University of Maine Roman Dwilewicz, University of Missouri–Rolla John J. Edgell, Jr., Southwest Texas State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Truman State University Russell Euler, Northwest Missouri State University Joe Flowers, San Antonio, Texas Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Tennessee Tech University Lou Grimm, University of Missouri–Rolla Raymond J. Grinnell, University of the West Indies Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Dennis R. Harmon, Missouri Southern State University Stan Hartzler, Virginia Polytech Institute & State University Troy Hicks, University of Missouri–Rolla Joe Howard, New Mexico Highlands University Timothy P. Keller, George Mason University Robert E. Kennedy, Pleasant Hill, Missouri Sava Krstic, Oregon Health & Science University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Mulatu Lemma, Savannah State University Edward S. Letzter, Temple University Man Chun Leung, National University of Singapore Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State University Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Phoebe McLaughlin, Central Missouri State University Peter R. Mercer, St. Mary’s College of Maryland Ilene Morgan, University of Missouri–Rolla

List of Referees (cont.)

Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Kent M. Neuerburg, Southeastern Louisiana University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Don Redmond, Southern Illinois University Les Reid, Southwest Missouri State University Jeffrey A. Rosoff, Gustavus Adolphus College Prasanna Sahoo, University of Louisville Jau–Shyong Shiue, University of Nevada–Las Vegas Donald P. Skow, University of Texas–Pan American Randall J. Swift, Western Kentucky University Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Mathew Timm, Bradley University Alvin Tinsley, Central Missouri State University Peter L. Vachuska, University of Wisconsin–Washington Co. Joseph Wiener, University of Texas–Pan American Bobby Neal Winters, Pittsburg State University Fred Worth, Henderson State University